Provide all details and eveluate the integral dx/e^x square root of 1-e^_2x
1
Expert's answer
2021-08-23T04:47:15-0400

Given integral is, "I = \int \frac{dx}{e^x\sqrt{1-e^{-2x}}}"


"I = \int \frac{dx}{e^x\sqrt{1-e^{-2x}}} = \int \frac{e^{-x}dx}{\sqrt{1-e^{-2x}}}"


Let "e^{-x} = t \implies e^{-x}dx = -dt"


"I= -\int \frac{dt}{\sqrt{1-t^2}}"


Using integration formula, "\int \frac{dx}{\sqrt{1-x^2}} = sin^{-1}(x) + C"


Then,


"I= -\int \frac{dt}{\sqrt{1-t^2}} =- sin^{-1}(t) + C"


Putting value of t,


"I = -sin^{-1}(e^{-x}) + C"